OpenStudy (anonymous):

using scientific notation find the logarithm of each number i dnt know how to do this at all two example log 12 and log 0.038 plus dnt know how to read the table feel so dumb

OpenStudy (anonymous):

ok, let's say you want to find log 12 (i'm assuming it's base 10?) we can write this as \[\log(1.2*10^1)=\log(1.2)+\log(10^1)=\log(1.2)+1\] that first step is an important rule with logarithms. Now, if you have a table, let's find 1.2 on it. It's probably broken down into sections based on the first digit or two, and then you have to read across. There's probably a 1.2 row, so you look in that row and go over to the 0 column; this means you're finding the log of 1.20, which is clearly the same as 1.2 since i'm not sure what your table is like, try it and tell me if that's how it's set up. If not, describe it to me (or link)

OpenStudy (anonymous):

thats how its set up

OpenStudy (anonymous):

sweet, then all you do is take that number and add 1 to it. similarly, for log 0.038, you put it in scientific notation, break it up like I did, and then find the value of 3.8 in the log table, then add.

OpenStudy (anonymous):

give me one sec

OpenStudy (anonymous):

so for 1.2 it would = 0.0792

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so do i move the decimal once or twice because you broke 0.038 into 3.8 and 12 into 1.2

OpenStudy (anonymous):

you move it as many times as necessary to make it so that there is one digit to the left of the decimal point. Then, the exponent on the ten becomes the number of places you moved it. If you move it to the left, the exponent, is positive, and if it's to the right, it's negative.

OpenStudy (anonymous):

for example, 34500 is 3.45*10^4 and .0065 is 6.5*10^(-3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about this one 21,100

OpenStudy (anonymous):

well, how many places would you have to move the decimal point to make it so that exactly 1 digit is on the left side of it?

OpenStudy (anonymous):

1 time

OpenStudy (anonymous):

2.1100

OpenStudy (anonymous):

how do i find that on the table

OpenStudy (anonymous):

that's 4, not 1... you go to the 2.1 row, then over to the 1 column

OpenStudy (anonymous):

owww your right forgot about that

OpenStudy (anonymous):

but since its 2.11 i find 2.1 the go over to one right

OpenStudy (anonymous):

then*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah buddy im getting the hang of this

OpenStudy (anonymous):

woo :D

OpenStudy (anonymous):

thing is im in the ACE program and the book doesnt really explain it as a one on one so its much harder

OpenStudy (anonymous):

well, i'm glad to be of help

OpenStudy (anonymous):

im glad to be helped lol